NicholasHen1615:

Algebra 2: What are the solutions of the system of equations y = –(x + 2)2 + 1 and y = 2x + 5? (–4, –3) and (–2, 1) (–4, –3) and (2, 1) (–4, 5) and (2, 1) (–4, 5) and (–2, 1)

OLIVER69:

You would need to subtract the equations first then simplify them Step One: y-y=-\left(x+2\right)^2+1-\left(2x+5\right) Step Two (simplify): 0 = -x^2 -6x - 8 Next you would find the solutions for the equation in step 2 which is x = -4 and -2 Then you would plug each one of the solutions in for the first equation or the second whichever you choose, I chose the first, so after I plugged them in I got the answer of (-4, -3) and (-2,1) (hope this helps)

NicholasHen1615:

OLIVER69:

Okay so for the first one I would start with the g(x) equation to help you find the correct answer first. I went through each equation and it can't be the first one because the g(x) equation is the red graph which means it should not be facing up And it's not the second or last one because the vertex of the graph should not be below the x-axis because the y-intercept is positive 3, so your answer is the third one.

NicholasHen1615:

Alright, thank you very much.